OpenStudy (anonymous):

f(x)=(x+3)(6-x)/(x-2)(x+3). State the the x-coordinate of any holes in the graph of this function. I got 0. Is that right? Please check my work. Thank you.

OpenStudy (anonymous):

Holes would occur whenever the denominator is zero... so when is the denominator (x-2)(x+3) zero?

OpenStudy (anonymous):

never?

OpenStudy (anonymous):

(x-2)=0 (x-3)=0

OpenStudy (anonymous):

Of course it can :) For a product of two numbers ab to be zero ab = 0 You only need ONE of them (either a or b) to be zero. Having said that (x-2)(x+3) = 0 You only need one of these two factors to be zero :)

OpenStudy (anonymous):

can I please get a visual?

OpenStudy (anonymous):

Is it 2?

OpenStudy (anonymous):

Okay, here's an example :) (x - 20) (x + 13) = 0 You only need one of these two factors to be zero, so either x - 20 = 0 OR x + 13 = 0 So what are the two possibilities?

OpenStudy (anonymous):

X=20 or x=-13

OpenStudy (anonymous):

That's right :) So, back to your denominator... When, then, is (x-2)(x+3) equal to zero?

OpenStudy (anonymous):

x=2 or x=-3

OpenStudy (anonymous):

And those... are when your function has holes ^.^

OpenStudy (anonymous):

Okay. So would the answer be 2?

OpenStudy (anonymous):

2, and -3 Remember, whenever x equals any of them (2 or -3) the denominator becomes zero, and the function cannot exist there, hence the hole.

OpenStudy (anonymous):

okay. i understand. thank u. ^_^

OpenStudy (anonymous):

No problem ^.^

OpenStudy (anonymous):

WAIT Are you still there?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I'm sorry, a hole is when BOTH the numerator and denominator are zero So you also have to find when the numerator (x+3)(6-x) =0 When is the numerator zero?

OpenStudy (anonymous):

when they both are different?

OpenStudy (anonymous):

No, do the same thing with the denominator, the way you found that it becomes zero when x=2 or x=-3 (x+3)(6-x) = 0

OpenStudy (anonymous):

x=-3 or x=6

OpenStudy (anonymous):

That's right So the numerator is zero when x = -3 or x = 6 The denominator is zero when x = 2 or x = -3 When are they BOTH zero

OpenStudy (anonymous):

i'm not sure.

OpenStudy (anonymous):

@PeterPan i came up with -3

OpenStudy (anonymous):

Yeah, that's correct :)

OpenStudy (anonymous):

Thank you. ^_^

OpenStudy (anonymous):

No problem :)